What is the free cash flow of a firm with revenues of $271 million, operating profit margin of 43%, tax rate of 25%, depreciation and amortization expense of $26 million, capital expenditures of $34 million, acquisition expenses of $8 million and change in net working capital of $14 million? Answer in millions, rounded to one decimal place (e.g., $245.63 = 245.6).
Answer #1

NOPAT = (1 - 0.25)[0.43(271) - 26] = $67.90 million

FCFF = 67.90 + 26 - 34 - 8 - 14

FCFF = $37.90 million